OpenStudy (nuccioreggie):

There are (32)4 ⋅ 30 bacteria in a petri dish. What is the total number of bacteria in the dish? 38 39 311 324

OpenStudy (nuccioreggie):

@DanJS

OpenStudy (nuccioreggie):

@Michele_Laino

Nnesha (nnesha):

use ^ for exponents like 2 to the 3rd power would be 2^3

Nnesha (nnesha):

\[\huge\rm (3^2)^4 \times 3^0\] like this ?

OpenStudy (nuccioreggie):

6561

Nnesha (nnesha):

anything to the zero power equal to one \[\huge\rm x^0 =1~~~~(Anything)^0=1\]

Nnesha (nnesha):

don't use calculator !

Nnesha (nnesha):

exponent rule \[\huge\rm (x^m)^n = x^{m \times n}\] just multiply the exponents base would stay the same

OpenStudy (nuccioreggie):

I GOT IT ITS a THANKS

OpenStudy (danjs):

sorry, just saw this ... i can't tell what the question actually is in order to get one of those answers

Nnesha (nnesha):

your answer choices are 3^8 3^9 3^(11) 3^(24)

OpenStudy (nuccioreggie):

its A BUT CAN U HELP ME WITH THIS ONE

OpenStudy (danjs):

oh, haha, right on sorry

OpenStudy (nuccioreggie):

What is the simplified expression for 2 power 2 multiplied by 2 power 3 over 2 power 4? 20 21 22 23

OpenStudy (danjs):

i tried all power combinations for the expression, but not the answers...

OpenStudy (danjs):

\[\huge\frac{ 2^2 * 2^3 }{ 2^4 }\]

OpenStudy (danjs):

since they all have the same base 2, and multiplied, you can combine that into a single term 2^ (power)

OpenStudy (danjs):

have to memorize and practice a little the like 4 general power rules...

OpenStudy (nuccioreggie):

SO ITS C?

OpenStudy (danjs):

http://www.rapidtables.com/math/number/exponent.htm scroll to the table, the first 5 riles is prolly good

OpenStudy (danjs):

rules

OpenStudy (danjs):

numerator, multiplying like bases, add the powers together

OpenStudy (danjs):

2^5 / 2 ^4

OpenStudy (danjs):

to move the bottom to the top, change the sign on the exponent, then use the addition of powers again

Nnesha (nnesha):

\[\huge\rm x^m \times x^n=x^{m+n}\] when we multiply same bases we should `ADD` their exponents and when you divide them you should `subtract` ther exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (danjs):

2^5 * 2^(-4)

OpenStudy (danjs):

= 2 or 2^1

OpenStudy (danjs):

haha, you have those latex things saved for all the basic things?

OpenStudy (nuccioreggie):

it B

OpenStudy (nuccioreggie):

A closed container has 3.06 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 0.47 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 ⋅ 10–1 0.51 grams, because (3.06 ⋅ 1.67) ⋅(1023 ⋅ 10–24) = 5.1102 ⋅ 10–1 4.73 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 5.11 grams, because (3.06 ⋅ 1.67) ⋅ (1023 ⋅ 10–24) = 5.1102

OpenStudy (danjs):

ok, you good on those rules now

OpenStudy (nuccioreggie):

i just need help with couple more lol

OpenStudy (nuccioreggie):

A closed container has 3.06 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 0.47 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 ⋅ 10–1 0.51 grams, because (3.06 ⋅ 1.67) ⋅(1023 ⋅ 10–24) = 5.1102 ⋅ 10–1 4.73 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10–24) = 4.73 5.11 grams, because (3.06 ⋅ 1.67) ⋅ (1023 ⋅ 10–24) = 5.1102

OpenStudy (danjs):

i-c